Sketch a line passing through the point (2, 1) and having
slope -1/2

Answer by stanbon(75887)

Sketch a line passing through the point (2, 1) and having
slope -1/2
--------------------
Form: y = mx+b
y = 1 when x = 2, and m = -1/2.
Solve for "b":
1 = (-1/2)2 + b
1 = -1 + b
b = 2
--------
Equation:
y = (-1/2)x+2
===================
Cheers,
Stan H.